Blast it! said David Wilson, president of Teledex Company. Weve just lost the bid on the Koopers job by $4,000. It seems
Answer #1
Required 1A :
Predetermined overhead rate = Total estimated manufacturing overhead / Total estimated direct labor = 898800 / 642000 140% of direct labor cost
Required 1B :
Manufacturing overhead cost applied = Actual total direct labor cost * Predetermined overhead rate = 13900 * 140% 19460
Required 2A :
Predetermined overhead rate ( Estimated Manufacturing overhead / Estimated Direct labor )
Fabricating department ( 374500 / 214000 ) 175% of direct labor cost
Machining department ( 428000 / 107000 ) 400% of direct labor cost
Assembly department ( 96300 / 321000 ) 30% of direct labor cost
Required 2B:
Manufacturing overhead cost applied = ( 5600 * 175% ) + ( 700 * 400% ) + ( 7600 * 30% ) 14880
Required 4A :
Koopers Job
Direct materials 7600
Direct labor 13900
Manufacturing overhead applied 19460
Total manufacturing cost 40960
Company's bid price ( @ 150% of total manufacturing cost ) 61440
Required 4B :
Koopers Job
Direct materials 7600
Direct labor 13900
Manufacturing overhead applied 14880
Total manufacturing cost 36380
Company's bid price ( @ 150% of total manufacturing cost ) 54570